# Avoiding Copyright Infringement via Large Language Model Unlearning

### v2

@@ 6.3 General-purpose Language Abilities
Figures 3(d), 6(d), and 7(d), show that GA, Gradient Difference, and TV both have sudden significant degradation of general-purpose abilities at certain time steps. NPO consistently underperforms SSU for Llama3.1 outperforms for Mistral-7B. Prompting and MemFree maintain stable general-purpose language abilities, close to the vanilla model.
@@ 7 Analysis
We analyze the impact of different components of SSU, including weight saliency maps and random labeling loss, on the sequential unlearning process. Figure 5 presents results for Llama3.1-8B-Instruct, and Figure 9 shows ablation results for Mistral-7B-Instruct-v0.3.
@@ 7.2 Impact of Random Labeling Loss
It is also noteworthy that, as shown in Figures 5(c) and 9(c), SSU without weight saliency and SSU without random loss lead to lower average Rouge scores for $D_{nor}$ in both models. This observation highlights the need for further research on the impact of unlearning algorithms on non-targeted knowledge retention, particularly in preserving content not intended for unlearning.
@@ 8 Conclusion
In this work, we explore the setting of unlearning copyrighted content sequentially from LLMs to avoid copyright infringement. We propose SSU, which utilizes random labeling loss and gradient-based weight saliency to achieve more stable sequential unlearning. Experiments demonstrate that SSU achieves a better trade-off between unlearning efficacy and general-purpose language abilities compared to existing methods. Future research could explore the effect of non-targeted knowledge redistribution during unlearning and provide theoretical certificate by connecting unlearning with Differential Privacy  Dwork (2006).
9 Limitations
In this work, we primarily use lexical-based evaluation metrics to evaluate the algorithm. However, as Ippolito et al. (2023) notes, measuring verbatim memorization might provide a false sense of privacy. Therefore, in addition to using Rouge-1 and Rouge-L scores, it is important to incorporate methods capable of detecting potential training data leakage. Membership Inference Attacks (MIAs) Shokri et al. (2017) offer a promising direction. Nonetheless, current research indicates that the performance of MIAs is near random guessing for pre-trained LLMs in various settings Duan et al. (2024); Yao et al. (2024). We encourage future research to develop more effective MIAs and apply them to our sequential unlearning setting.
Furthermore, although SSU achieves a better trade-off between unlearning efficacy and general-purpose language abilities retention compared to other baseline methods, we still observe some unintended knowledge loss in books that are not targeted for unlearning, as well as a reduction in the model’s reasoning and conversation abilities. Future work should focus on further minimizing the gap in knowledge and capability retention between the modified model and the original, in order to ensure better general knowledge preservation during sequential unlearning.
Lastly, our work does not provide a theoretical guarantee to certify unlearning. Recently, Zhang et al. (2024a) explored certified unlearning for deep neural networks in non-convex settings by establishing a connection between unlearning and differential privacy Dwork (2006). A similar certified mechanism is also necessary for transformer-based generative models, and we leave this as an avenue for future work.
Appendix A Appendix: Experiment Details
@@ A.2.1 Rouge-1
$Recall=\frac{\text{overlapping unigrams}}{\text{total unigrams in the reference text}}.$ (9)
@@
$Precision=\frac{\text{overlapping unigrams}}{\text{total unigrams in the hypothesis text}}.$ (10)
@@
$F1=2\cdot\frac{Precision\cdot Recall}{Precision+Recall}$ (11)
@@ A.2.2 Rouge-L
$Recall=\frac{LCS}{\text{length of the reference text}}.$ (12)
@@
$Precision=\frac{LCS}{\text{length of the hypothesis text}}.$ (13)
@@
$F1=2\cdot\frac{Precision\cdot Recall}{Precision+Recall}$ (14)
@@ A.4.3 Unlearning via Gradient Difference
$\mathcal{L}_{\text{fgt}}=-\sum_{(x_{\text{fgt}},y_{\text{fgt}})\in D_{f}}L(x_{\text{fgt}},y_{\text{fgt}},\theta_{t})$ (16)
$\mathcal{L}_{\text{rnd}}:=\sum_{(x_{\text{fgt}},)\in D_{f}}\frac{1}{|Y_{\text{rnd}}|}\sum_{(,y_{\text{rnd})}\in Y_{\text{rnd}}}L(x_{\text{fgt}},y_{\text{rnd}},\theta_{t})$ (17)
$\phi_{\theta}=h_{\theta}(x_{\text{nor}},y_{\text{nor}<i})$ (18)
$\mathcal{L}_{\text{add}}:=\sum_{(x_{\text{add}},y_{\text{add}})\in D_{\text{add}}}\sum_{i=1}^{|y_{\text{add}}|}\text{KL}(\phi_{\theta_{o}}\parallel\phi_{\theta_{t}}).$ (19)
@@
$L=\epsilon_{1}\mathcal{L}_{\text{fgt}}+\epsilon_{2}\mathcal{L}_{\text{rnd}}+\epsilon_{3}\mathcal{L}_{\text{add}}$ (20)
@@ A.4.4 Unlearning via Task Vector
$\mathcal{L}_{\text{fgt}}=\sum_{(x_{\text{fgt}},y_{\text{fgt}})\in D_{f}}L(x_{\text{fgt}},y_{\text{fgt}},\theta_{t})$ (21)
$\theta_{t+1}\leftarrow\theta_{t}-\epsilon\nabla_{\theta_{t}}\mathcal{L}_{\text{fgt}}$ (22)
@@ A.4.5 Unlearning via NPO
$\mathcal{L}_{\text{NPO}}=\frac{2}{\beta}\mathbb{E}_{(x,y)\in D_{\text{forget}}}\left[\log\left(1+\left(\frac{\pi_{\theta}(y|x)}{\pi_{\text{ref}}(y|x)}\right)^{\beta}\right)\right]$ (25)
@@ A.5 Implementation Details
For the Llama-3.1-8B-Instruct model, we set the learning rate to be 1e-5 for the first five time steps, and decrease the learning rate to 1e-6 for the rest of the time steps.
For the Mistral-7B-Instruct-v0.3, we set the learning rate to be 1e-5 for the first three time steps, and the learning rate to be 1e-6 for the rest of time steps.
@@
SSU. For SSU, we set $\epsilon_{1}=1$,a and $\epsilon_{2}=0.5$ for all the time steps and models. We set $\gamma$ to be 1 standard deviation away from the mean of the gradient vector $\nabla_{\theta}L_{f}(\theta_{t})$
Appendix B Appendix: Full Experiment Results
@@
Unlearning $\text{Unlearning Efficacy}=U_{f}+U_{\text{prev}}$ (26)
@@
and
$U_{\text{prev}}=-\frac{1}{2}\left(R_{1}^{\text{prev}}+R_{L}^{\text{prev}}\right)$
Here, $R_{1}^{f}$ and $R_{L}^{f}$ represent the Rouge-1 and Rouge-L scores for $D_{f}$, and $R_{1}^{\text{prev}}$ and $R_{L}^{\text{prev}}$ represent the Rouge-1 and Rouge-L scores for $D_{prev}$.
The y-axis is defined as:
[…]
@@
Appendix C Appendix: Numeric Experiment Results
In this section, we present our experimental results numerically. Tables 1 to 10 display the unlearning results of Llama3.1 across all ten time steps. Tables 11 to 16 show the results for Mistral-7B unlearning across all six time steps. In sections C.1 and C.2 we illustrate how GA, TV, and Gradient Difference encounter catastrophic collapse.
C.1 Catastrophic Collapse of Llama3.1
GA, TV, and Gradient Difference experience varying levels of model collapse during the sequential unlearning process. As shown in table 1, GA begins with an MMLU of 0.5821. By time step 5, GA’s MMLU has dropped to 0.3102 5, demonstrating a rapid degradation in general reasoning. Similarly, TV starts with an MMLU of 0.6621 at time step 1 and undergoes a steep decline in reasoning ability, dropping to 0.4887 by time step 5, and reaching an MMLU of 0 at time step 10 (as shown in tables 5 and 10). Additionally, Gradient Difference also faces catastrophic collapse at time step 5. Specifically, its MT-Bench score falls from 8.034 at time step 4 (table 4) to 4.9438 at time step 5, and further declines to 4.48 at time step 10, though its MMLU score remains stable. Lastly, NPO and SSU exhibit a gradual decline in general-purpose language abilities, but SSU consistently outperforms NPO across all the time steps.
C.2 Catastrophic Collapse of Mistral-7B
GA and TV experience rapid model collapse, while Gradient Difference still suffers loss of conversational ability. As shown in table 11, GA collapses at the first time step. TV initially declines gradually but undergoes a sudden reasoning degradation at time step 5 (table 14), where both MMLU and MT-Bench scores drop to 0. Gradient Difference experiences a sharp decrease in MT-Bench score from 7.2375 at time step 4 to 2.8375 at time step 5, eventually dropping to 2.6815 at time step 6. Similar to the Llama3.1 case, Gradient Difference maintains stable performance on MMLU. Lastly, NPO maintains competitive general-purpose abilities scores compared to SSU, Figures 7 and 8 show that SSU demonstrates superior unlearning efficacy and achieves a better trade-off.

### v3

@@ 6.3 General-purpose Language Abilities
Figures 3(d), 6(d), and 9(d), show that GA, Gradient Difference, and TV both have sudden significant degradation of general-purpose abilities at certain time steps. NPO consistently underperforms SSU for Llama3.1 outperforms for Mistral-7B. Prompting and MemFree maintain stable general-purpose language abilities, close to the vanilla model.
@@ 7 Analysis
We analyze the impact of different components of SSU, including weight saliency maps and random labeling loss, on the sequential unlearning process. Figure 5 presents results for Llama3.1-8B-Instruct, and Figure 11 shows ablation results for Mistral-7B-Instruct-v0.3. We further present the effect of varying $\epsilon_{1}$ and $\epsilon_{2}$ in equation 4 in appendix B.1.
@@ 7.2 Impact of Random Labeling Loss
It is also noteworthy that, as shown in Figures 5(c) and 11(c), SSU without weight saliency and SSU without random loss lead to lower average Rouge scores for $D_{nor}$ in both models. This observation highlights the need for further research on the impact of unlearning algorithms on preserving content not intended for unlearning.
@@ 8 Conclusion
We explore sequential unlearning of copyrighted content in LLMs to mitigate copyright infringement. We propose SSU, which utilizes random labeling loss and gradient-based weight saliency to achieve more stable sequential unlearning. Experiments show SSU outperforms existing methods in balancing unlearning efficacy and language retention. Future research could focus on enhancing the robustness of unlearning algorithms and exploring other copyright takedown methods to further mitigate the risks of copyright infringement.
Limitations
Robustness of Machine Unlearning
Our evaluation relies on lexical-based metrics, which, as noted by Ippolito et al. (2023), may provide a false sense of privacy. Additionally, Łucki et al. (2024) and Cooper et al. (2024) highlight that existing unlearning methods often lack robustness and may merely obscure data rather than true unlearning.
In our experiments, we observe that knowledge from previously unlearned books can re-emerge at certain future time steps, a phenomenon present across all unlearning algorithms tested (see Appendix B.2). This further underscores the challenges of sequential unlearning and the necessity for more robust unlearning algorithms.
Unintended Knowledge and Capability Loss
While SSU achieves a better trade-off between unlearning efficacy and general-purpose language retention, it still weakens the model’s reasoning and conversational abilities. Future work should further reduce this gap, ensuring better preservation of general knowledge during sequential unlearning.
Unlearning Is Not All You Need
Mitigating copyright infringement risks requires more than unlearning alone. For example, Liu et al. (2024a) introduces Infini-gram, an efficient search engine for scanning large-scale LLM training corpora. A complementary approach could involve using such tools to detect copyrighted passages in generated outputs and intervene to ensure non-copyrighted regeneration. This generation-time copyright takedown method represents a promising research direction. Future work could expand search engine tools to cover a broader range of training datasets used in state-of-the-art LLMs. Additionally, regularly updating model checkpoints to align with the right to be forgotten would be crucial for maintaining legal and ethical compliance.
Acknowledgements
We thank Bryan Li, Yikai Mao, and anonymous reviewers for their valuable feedback and insightful discussions on this paper.
This research is supported in part by the Office of the Director of National Intelligence (ODNI), the Intelligence Advanced Research Projects Activity (IARPA) via the BENGAL Program, the Defense Advanced Research Projects Activity (DARPA) via the SciFy Program, and Schmidt Sciences via the AI2050 fellowship. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of ODNI, IARPA, DARPA, Schmidt Sciences, or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for governmental purposes notwithstanding any copyright annotation therein.
Appendix A Experiment Details
@@ A.2.1 Rouge-1
$Recall=\frac{\text{overlapping unigrams}}{\text{total unigrams in the reference}}.$
@@
$Precision=\frac{\text{overlapping unigrams}}{\text{total unigrams in the hypothesis}}.$
@@
$F1=2\cdot\frac{Precision\cdot Recall}{Precision+Recall}$
@@ A.2.2 Rouge-L
$Recall=\frac{LCS}{\text{length of the reference text}}.$
@@
$Precision=\frac{LCS}{\text{length of the hypothesis text}}.$
@@
$F1=2\cdot\frac{Precision\cdot Recall}{Precision+Recall}$
@@ A.4.3 Unlearning via Gradient Difference
$\mathcal{L}_{\text{fgt}}=-\sum_{(x_{\text{fgt}},y_{\text{fgt}})\in D_{f}}L(x_{\text{fgt}},y_{\text{fgt}},\theta_{t})$
$\mathcal{L}_{\text{rnd}}:=\sum_{(x_{\text{fgt}},)\in D_{f}}\frac{1}{|Y_{\text{rnd}}|}\sum_{(,y_{\text{rnd})}\in Y_{\text{rnd}}}L(x_{\text{fgt}},y_{\text{rnd}},\theta_{t})$
$\phi_{\theta}=h_{\theta}(x_{\text{nor}},y_{\text{nor}<i})$
$\mathcal{L}_{\text{add}}:=\sum_{(x_{\text{add}},y_{\text{add}})\in D_{\text{add}}}\sum_{i=1}^{|y_{\text{add}}|}\text{KL}(\phi_{\theta_{o}}\parallel\phi_{\theta_{t}}).$
@@
$L=\epsilon_{1}\mathcal{L}_{\text{fgt}}+\epsilon_{2}\mathcal{L}_{\text{rnd}}+\epsilon_{3}\mathcal{L}_{\text{add}}$
@@ A.4.4 Unlearning via Task Vector
$\mathcal{L}_{\text{fgt}}=\sum_{(x_{\text{fgt}},y_{\text{fgt}})\in D_{f}}L(x_{\text{fgt}},y_{\text{fgt}},\theta_{t})$
$\theta_{t+1}\leftarrow\theta_{t}-\epsilon\nabla_{\theta_{t}}\mathcal{L}_{\text{fgt}}$
@@ A.4.5 Unlearning via NPO
$\mathcal{L}_{\text{NPO}}=\frac{2}{\beta}\mathbb{E}_{(x,y)\in D_{\text{f}}}\left[\log\left(1+\left(\frac{\pi_{\theta}(y|x)}{\pi_{\text{ref}}(y|x)}\right)^{\beta}\right)\right]$
@@ A.5 Implementation Details
For the Llama-3.1-8B-Instruct model, we set the learning rate to be 1e-5 for the first five time steps, and decrease the learning rate to 1e-6 for the rest of the time steps. For the Mistral-7B-Instruct-v0.3, we set the learning rate to be 1e-5 for the first three time steps, and the learning rate to be 1e-6 for other time steps. For SSU, we set $\epsilon_{1}=1$,a and $\epsilon_{2}=0.5$ for all the time steps and models. We set $\gamma$ to be 1 standard deviation away from the mean of the gradient vector $\nabla_{\theta}L_{f}(\theta_{t})$
Appendix B Additional Analysis
B.1 Impact of $\epsilon_{1}$ and $\epsilon_{2}$
In this section, we analyze the impact of the hyperparameters $\epsilon_{1}$ and $\epsilon_{2}$ in Equation 4. Specifically, we denote $SSU_{1.5,1}$ as the method with $\epsilon_{1}=1.5$ and $\epsilon_{2}=1$ on Llama 3.1, and $SSU_{1,2}$ as the method with $\epsilon_{1}=1$ and $\epsilon_{2}=2$. All other hyperparameters remain consistent with the settings used in our main experiments.
As shown in Table 1, SSU maintains strong performance on MMLU and MT-Bench, indicating its stability across different parameter values. Unlike gradient ascent and gradient difference methods, SSU does not lead to model collapse. However, in our main experiments, we tuned $\epsilon_{1}$ and $\epsilon_{2}$ to achieve the best trade-off between unlearning efficacy and general-purpose language abilities.
Time Step Method $D_{f}$ $D_{f}$ $D_{prev}$ $D_{prev}$ $D_{nor}$ $D_{nor}$ MMLU MT-Bench
1 $SSU_{1.5,1}$ 0.2513 0.1424 N/A N/A 0.2307 0.1335 0.6631 8.1180
1 $SSU_{1,2}$ 0.2515 0.1417 N/A N/A 0.2316 0.1335 0.6624 7.8058
2 $SSU_{1.5,1}$ 0.2856 0.1688 0.2558 0.1433 0.2402 0.1376 0.6463 8.1850
2 $SSU_{1,2}$ 0.2852 0.1689 0.2567 0.1433 0.2393 0.1367 0.6470 8.1675
3 $SSU_{1.5,1}$ 0.2415 0.1400 0.2699 0.1539 0.2403 0.1378 0.6463 7.9870
3 $SSU_{1,2}$ 0.2414 0.1386 0.2704 0.1543 0.2394 0.1370 0.6466 8.1283
4 $SSU_{1.5,1}$ 0.2492 0.1390 0.2576 0.1517 0.2458 0.1389 0.6431 8.0349
4 $SSU_{1,2}$ 0.2420 0.1403 0.2565 0.1511 0.2439 0.1391 0.6431 8.0266
5 $SSU_{1.5,1}$ 0.2551 0.1365 0.2650 0.1527 0.2441 0.1385 0.6298 7.9812
5 $SSU_{1,2}$ 0.2545 0.1369 0.2640 0.1517 0.2429 0.1378 0.6319 8.0578
6 $SSU_{1.5,1}$ 0.2494 0.1412 0.2621 0.1472 0.2435 0.1392 0.6298 7.9812
6 $SSU_{1,2}$ 0.2476 0.1408 0.2632 0.1467 0.2433 0.1380 0.6263 8.0484
Table 1: Results across time steps with different $\epsilon$. All performance metrics on $D$ are reported as Rouge-L scores.
B.2 Re-Emergence in Sequential Unlearning
In the main text, we use $D_{prev}$the set of all previously unlearned books—to assess whether they remain unlearned throughout sequential unlearning. Here, instead of evaluating all unlearned books collectively, we examine how knowledge of each previously unlearned book evolves as new books are unlearned at later time steps.
Figures 7 and 8 present the results for Llama 3.1 and Mistral-7B, respectively. The $x$-axis denotes the number of unlearning steps performed, while the $y$-axis represents the ROUGE-L score for each individual book. Notably, in certain future time steps, the ROUGE-L score of previously unlearned books increases. For instance, as shown in Figure 7(d), the ROUGE-L score of the fourth book under NPO rises as the model unlearns additional books. Similar trends appear in Figure 7(a), where NPO’s score for the first book increases at time step 9, and in Figure 8(a), where the Mistral model exhibits a steady increase in NPO’s score as it unlearns other books over time.
We also observe instability in the gradient difference, particularly in Figure 7, where performance on previously unlearned books fluctuates significantly across future time steps. Additionally, SSU exhibits some degree of reappearance of previously unlearned knowledge. This aligns with Łucki et al. (2024), in which it pointed out that fine-tuning on data unrelated to $D_{f}$ can inadvertently recover unlearned knowledge. Since all baselines in our experiments rely on fine-tuning with unlearning objectives, sequential unlearning fine-tunes the model on unrelated data—books unlearned at later time steps are independent of those unlearned earlier. This highlights a key challenge in sequential unlearning: continual model modifications may lead to knowledge re-emergence, highlighting the need for more robust unlearning algorithms in sequential settings.
(a) ROUGE-L Score for Book 1
(b) ROUGE-L Score for Book 2
(c) ROUGE-L Score for Book 3
(d) ROUGE-L Score for Book 4
(e) ROUGE-L Score for Book 5
(f) ROUGE-L Score for Book 6
(g) ROUGE-L Score for Book 7
(h) ROUGE-L Score for Book 8
Figure 7: ROUGE-L score for each previously unlearned book, evaluated using various unlearning methods on the Llama 3.1. The x-axis represents the number of time steps (i.e., the number of sequential unlearning operations performed), while the y-axis indicates the ROUGE-L score. Different unlearning methods are compared, with ‘Vanilla’ serving as the baseline model without unlearning. We omit certain time steps for some methods because of the catastrophic collapse.
(a) ROUGE-L Score for Book 1
(b) ROUGE-L Score for Book 2
(c) ROUGE-L Score for Book 3
(d) ROUGE-L Score for Book 4
(e) ROUGE-L Score for Book 5
(f) ROUGE-L Score for Book 6
Figure 8: ROUGE-L score for each previously unlearned book, evaluated using various unlearning methods on the Mistral model. The x-axis represents the number of time steps (i.e., the number of sequential unlearning operations performed), while the y-axis indicates the ROUGE-L score. Different unlearning methods are compared, with ‘Vanilla’ serving as the baseline model without unlearning. We omit certain time steps for some methods because of the catastrophic collapse.
Appendix C Full Experiment Results
@@
Unlearning $\text{Unlearning Efficacy}=U_{f}+U_{\text{prev}}$
@@
$U_{\text{prev}}=-\frac{1}{2}\left(R_{1}^{\text{prev}}+R_{L}^{\text{prev}}\right).$
Here, $R_{1}^{f}$ and $R_{L}^{f}$ represent the Rouge-1 and Rouge-L scores for $D_{f}$, and $R_{1}^{\text{prev}}$ and $R_{L}^{\text{prev}}$ represent the Rouge-1 and Rouge-L scores for $D_{prev}$. The y-axis (General-purpose Language Ability) is defined as:
$\frac{1}{2}\left(M+\frac{B}{10}\right)$
@@
Appendix D Numeric Experiment Results
In this section, we present our experimental results numerically. Tables 2 to 11 display the unlearning results of Llama3.1 across all ten time steps. Tables 12 to 17 show the results for Mistral-7B unlearning across all six time steps. In sections D.1 and D.2 we illustrate how GA, TV, and Gradient Difference encounter catastrophic collapse.
D.1 Catastrophic Collapse of Llama3.1
GA, TV, and Gradient Difference experience varying levels of model collapse during the sequential unlearning process. As shown in table 2, GA begins with an MMLU of 0.5821. By time step 5, GA’s MMLU has dropped to 0.3102 6, demonstrating a rapid degradation in general reasoning. Similarly, TV starts with an MMLU of 0.6621 at time step 1 and undergoes a steep decline in reasoning ability, dropping to 0.4887 by time step 5, and reaching an MMLU of 0 at time step 10 (as shown in tables 6 and 11). Additionally, Gradient Difference also faces catastrophic collapse at time step 5. Specifically, its MT-Bench score falls from 8.034 at time step 4 (table 5) to 4.9438 at time step 5, and further declines to 4.48 at time step 10, though its MMLU score remains stable. Lastly, NPO and SSU exhibit a gradual decline in general-purpose language abilities, but SSU consistently outperforms NPO across all the time steps.
D.2 Catastrophic Collapse of Mistral-7B
GA and TV experience rapid model collapse, while Gradient Difference still suffers loss of conversational ability. As shown in table 12, GA collapses at the first time step. TV initially declines gradually but undergoes a sudden reasoning degradation at time step 5 (table 15), where both MMLU and MT-Bench scores drop to 0. Gradient Difference experiences a sharp decrease in MT-Bench score from 7.2375 at time step 4 to 2.8375 at time step 5, eventually dropping to 2.6815 at time step 6. Similar to the Llama3.1 case, Gradient Difference maintains stable performance on MMLU. Lastly, NPO maintains competitive general-purpose abilities scores compared to SSU, Figures 9 and 10 show that SSU demonstrates superior unlearning efficacy and achieves a better trade-off.